OpenStudy (anonymous):

log x - log ( x + 3 ) = 1

OpenStudy (ranga):

Also, if \[\Large \log_{10} A = B \]then\[\Large A = 10^{B}\]

OpenStudy (ranga):

\[\Large \log(A) - \log(B) = \log(\frac{ A }{ B })\]

OpenStudy (ranga):

log x - log ( x + 3 ) = log( x/(x + 3) ) = 1 So x/(x + 3) = 10^1 = 10 Solve for x.